OpenStudy (anonymous):

Find the pH level, to the nearest tenth, of a liquid with [H+] about 6.5*10^-3 a.-3.8 b.3.8 c.2.2 d.3.0

jimthompson5910 (jim_thompson5910):

pH = -log(H+) pH = -log(6.5*10^(-3)) pH = -(-2.187) pH= 2.187 pH = 2.2 So the pH is 2.2, which makes the answer choice C

OpenStudy (anonymous):

thanks for working it out.

jimthompson5910 (jim_thompson5910):

sure thing